solve the following question:
||x^2-2x-8|+x^4-x^3+4x+8| =6
1
Expert's answer
2019-11-12T11:35:16-0500

Solution:

"|x^2-2\times x-8|+x^4-x^3+4\times x+8=6 or |x^2-2\times x-8|+x^4-x^3+4\times x+8=-6"

We construct graphs of the function where the polynomial obtained as a result of the transformation of the original equation is contained in the right-hand side. The equation will have solutions if there are intersection points of these graphs with the abscissa axis.

"|x^2-2\times x-8|+x^4-x^3+4\times x+8=6"

if "x^2-2\times x-8\ge0"

The resulting equation has no roots.

"x^4-x^3+x^2+2\times x-6=0"

illustrate this on the graphs



minimum 18 if

       x \leq-2    Does not cross the abscissa axis.


minimum 110 Does not cross the abscissa axis.

"x^2-2\times x-8\leq0"

we get the equation

"x^4-x^3+6\times x+10=0, -2<x<4"



minimum 6ю Does not cross the abscissa axis.

"|x^2-2\times x-8|+x^4-x^3+4\times x+8=-6"

if "x^2-2\times x-8\ge0" we will have eqation

"x^4-x^3+x^2+2\times x+6=0"


minimum 30 if   /Does not cross the abscissa axis.


minimum 222 if "x \ge 4"

if "x^2-2\times x-8\leq0" . we have eqation

"x^4-x^3-x^2+6\times x+22=0, -2<x<4"



minimum 17.

Since no intersection points of the constructed graphs with the abscissa were obtained, the equation has no roots.


Answer: The equation has no roots.